The following information was available for the year ended December 31, 2019:

Sales $ 500,000
Net income 71,700
Average total assets 800,000
Average total stockholders' equity 375,000
Dividends per share 1.35
Earnings per share 3.00
Market price per share at year-end 28.20


Required:

  1. Calculate margin, turnover, and ROI for the year ended December 31, 2019.
  2. Calculate ROE for the year ended December 31, 2019.
  3. Calculate the price/earnings ratio for 2019.
  4. Calculate the dividend payout ratio for 2019.
  5. Calculate the dividend yield for 2019.
Answer #1

a) Margin = Net income/Sales = 71700/500000 = 14.34%

Turnover = Sales/Average operating assets = 500000/800000 = 0.625

ROI = Net income/Average operating assets = 71700/800000 = 8.96%

b) ROE = 71700/375000 = 19.12%

c) Price earning ratio = MPS/EPS = 28.2/3 = 9.4

D) Dividend payout ratio = DPS/EPS = 1.35/3 = 45%

e) Dividend yield = DPS/MPS = 1.35/28.2 = 4.79%